A force in the +x direction with a magnitude of 18.0 N is applied to a 6.00 kg box that is sitting on the horizontal, frictionless surface of a frozen lake. If this is the only horizontal force on the box and it is initially at rest at x=0, what is the speed after it has traveled 14.0 m?
Answer #1

Mass of the box m = 6 Kg

force applied F = 18 N

accelaration in the box a = F/m = 18/6 = 3 m/s2

frome the relation a = ( v2 - u2 ) /2s

=  v2  /2s since u = 0

hence speed v = sq.root of 2as

= sqroot of (2 x 3 x14)

= 9.16 m/s